# ADAM-SINDy: An Efficient Optimization Framework for Parameterized Nonlinear Dynamical System Identification

### v2

@@ 1 Introduction
The paper is structured as follows. Section 2 provides an overview of the classical SINDy methodology, outlining its mathematical foundation and then introducing the ADAM-SINDy methodology, emphasizing its advantages and the implementation of the gradient descent algorithm. This section also details the master candidate library of nonlinear functions used in evaluating the classical SINDy and ADAM-SINDy frameworks. Section 3 examines benchmark problems across various applications, including ordinary and partial differential equations, focusing on problems that have nonlinear parameters, thereby challenging the classical SINDy method. Section 4 discusses the advantages of the ADAM-SINDy method over the classical approach. Finally, Section 5 presents a conclusion.
@@ 2 Methods
2.1 Sparse Identification of Nonlinear Dynamics (SINDy)
@@
where $\mathbf{x}(t)\in\mathbb{R}^{n}$ denotes the state vector at time $t$, $\dot{\mathbf{x}}(t)$ represents its time derivative, and $\mathbf{f}(\mathbf{x})$:::: $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a nonlinear function governing the system’s dynamics. The objective of the SINDy algorithm is to identify the function $\mathbf{f}(\mathbf{x}(t))$ from available data. The process begins by collecting a dataset consisting of measurements of $m$ states at $n$ discrete time-steps. The data is organized as follows:
@@ 2.2 ADAM Optimization Augmented SINDy: ADAM-SINDy
The ADAM-SINDy methodology, a novel extension of the traditional SINDy framework [34], introduces substantial improvements by adopting ADAM optimization in place of the more commonly employed iterative least squares algorithm for solving the LASSO problem. This methodological innovation significantly addresses key challenges in parameter estimation, while concurrently reducing the dependency on the predefined library of candidate functions, a key limitation that has often constrained the SINDy approaches. Unlike the classical SINDy framework, which heavily relies on the initial selection of candidate parameters, ADAM-SINDy adaptively identifies relevant terms from the candidate library $(\bm{\Xi})$ while simultaneously optimizing nonlinear parameters $(\bm{\Lambda})$. This adaptive process minimizes reliance on prior assumptions and enhances the capability to discover complex, nonlinearly defined parameters in a more robust manner. Nonlinear parameters, such as frequencies in trigonometric functions, bandwidths in exponential functions (e.g., kernels), and non-integer polynomial exponents, present challenges in the traditional SINDy approach, which requires prior knowledge of these parameters. ADAM-SINDy addresses this limitation by making these parameters optimizable within a global optimization framework. Mathematically, the sparse regression, Eq. 6, is then reformulated as
$\min_{\bm{\Xi}}\max_{\begin{subarray}{c}\lambda,\bm{\Gamma}\end{subarray}}\left[\|\dot{\mathbf{X}}-\bm{\Theta}(\mathbf{X}\textbf{;}\thinspace\bm{\Lambda})\thinspace\bm{\Xi}\|_{2}^{2}+\lambda\|\bm{\Gamma}\thinspace\bm{\Xi}\|_{1}\right]\;,$ (8)
@@
where the objective is minimized with respect to the coefficient matrix and maximized with respect to the sparsity knobs, respectively. This min-max approach is similar to the adaptive weighting approach in physics-informed neural networks where the weights for different losses is adaptively updated [57]. Moroever, in practice, either $\lambda$ or $\bm{\Gamma}$ is selected to enforce sparsity. Among these, $\bm{\Gamma}$ is preferred due to its ability to treat each candidate function individually. Iteratively updating the weighting matrix $\bm{\Gamma}$ shares similar strategies with the Iteratively Reweighted Least Squares (IRLS) technique, which is applied to both constrained [48] and unconstrained problems [50]. In the present work, we adopt a similar strategy by iteratively updating the $\bm{\Gamma}$ matrix, but with the distinction of updating the weights using gradients computed through backpropagation.
@@
    Adaptive weighting: ADAM-SINDy treats the regularization parameter, $\bm{\Gamma}$, as an optimizable matrix (inspired by the IRLS methodology), where the candidate-specific weights are also optimized over epochs. As a result, the method facilitates an individualized treatment of the candidate functions, thereby enhancing the accuracy of parameter identification and improving the overall robustness of the system identification process.
@@
The ADAM-SINDy framework, detailed in Algorithm 1, introduces a fully adaptive global optimization process for all parameters typically preselected in the classical SINDy framework. It begins with the initialization of key parameters—$\bm{\Xi}$, $\bm{\Lambda}$, $\lambda$, and $\bm{\Gamma}$—which are iteratively optimized based on the provided input data. To enhance computational efficiency and robustness, the input data is divided into batches of size $B$. When the batch size is set equal to the size of the input data, the algorithm operates under a full-batch optimization procedure. During each epoch, for each batch (or the full dataset in the case of full-batch optimization), the loss function $\mathcal{L}_{b}$ is computed to measure the discrepancy between the observed system dynamics, $\dot{\mathbf{X}}_{b}$, and the model’s predicted dynamics, represented by $\bm{\Theta}(\mathbf{X}_{b};\bm{\Lambda})\thinspace\bm{\Xi}$. This batch-wise loss computation incorporates a sparsity-inducing term, regulated by either $\lambda$ or $\bm{\Gamma}$ knob, ensuring that only the most relevant terms are selected.
The parameters—$\bm{\Xi}$, $\bm{\Lambda}$, $\lambda$, and $\bm{\Gamma}$—are then updated via the stochastic gradient descent algorithm on a per-batch basis, gradually minimizing the loss function across the entire dataset. In the case of full-batch optimization, updates are performed after computing the loss over the entire input dataset at once. A sequential thresholding step is employed after each batch update, where values of both $\bm{\Xi}$ and $\bm{\Lambda}$ that are small are set to zero if they fall below the specified tolerance, $\epsilon$. This iterative process, which repeats for each batch and continues until the maximum number of iterations, $N_{\text{max}}$, is reached, ensures that the optimization converges efficiently.
Algorithm 1 ADAM-SINDy Algorithm
Input:
 Observational data $\mathbf{X}$ and $\dot{\mathbf{X}}$
 Tolerance $\epsilon$ and maximum iterations $N_{\text{max}}$
Initialization:
 Set iteration counter $n\leftarrow 0$
Optimization Loop:
while $n<N_{\text{max}}$ do
      For each batch $b=1,2,\dots$
       Sample a batch of data $\mathbf{X}_{b}$ and $\dot{\mathbf{X}}_{b}$
       Compute $\bm{\Theta}(\mathbf{X}_{b};\bm{\Lambda})$
       Compute Loss Function:
$\mathcal{L}_{b}\leftarrow\left\|\dot{\mathbf{X}}_{b}-\bm{\Theta}(\mathbf{X}_{b};\bm{\Lambda})\bm{\Xi}\right\|_{2}^{2}+\lambda\left\|\bm{\Gamma}\bm{\Xi}\right\|_{1}$
       Update Parameters using Backpropagation:
        if Uniform Sparsity then
$\lambda_{n+1}\leftarrow\lambda_{n}+\eta_{\lambda}\nabla_{\lambda}\mathcal{L}_{b}(\bm{\Theta}_{b},\bm{\Xi},\bm{\Lambda},\lambda,\bm{\Gamma})$
        else
$\bm{\Gamma}_{n+1}\leftarrow\bm{\Gamma}_{n}+\eta_{\bm{\Gamma}}\nabla_{\bm{\Gamma}}\mathcal{L}_{b}(\bm{\Theta}_{b},\bm{\Xi},\bm{\Lambda},\lambda,\bm{\Gamma})$
      Increment iteration counter: $n\leftarrow n+1$
end while
Output:
 Optimized parameters $\bm{\Xi}^{*}$ and $\bm{\Lambda}^{*}$
2.2.1 Stochastic versus Full-Batch Optimization
The ADAM-SINDy framework provides the flexibility to implement both a batching strategy through stochastic gradient descent and a full-batch optimization approach. In the stochastic case, optimization is performed on smaller, randomly sampled subsets of the overall dataset. This method optimizes computational resource utilization by partitioning the data into manageable chunks, thereby reducing memory requirements and accelerating the optimization process. The batch size must be sufficiently large to capture the dynamics of interest while remaining small enough to facilitate efficient computation.
In contrast, full-batch optimization processes the entire dataset in a single iteration. This approach leverages all available information, potentially yielding more accurate and stable updates to the optimization parameters. Notably, ADAM optimization employs momentum, which can help avoid local minima even in the absence of stochasticity introduced by mini-batches.
@@ 2.3 Problem Statement
The proposed ADAM-SINDy methodology also offers further significant advantages due to the trainable nature of the nonlinear parameters. Specifically, when the parameters associated with cosine and exponential functions—$\mathcal{C}$ and $\mathcal{D}$—approach zero, their non-zero coefficients effectively interact with the constant term $1$ present in the library. Similarly, in the context of nonlinear polynomial combined terms, specifically those involving exponential and cosine functions, when $\mathcal{F}$ and $\mathcal{G}$ converge to zero, these terms interact with the polynomial terms $\mathbf{X}^{\mathcal{A}}$. This phenomenon of effective self-adaptation of nonlinear functions and its interaction with other functions, as distinctly observed in the ADAM-SINDy framework is particularly crucial in scenarios where predefined candidate functions lack an exact function required to accurately represent the true system dynamics, as the trainable parameters are optimized to compensate for these deficiencies.
@@ 3 Numerical Experiments
The results from ADAM-SINDy using batch-wise optimization are excluded, as this method failed to produce accurate outcomes for complex dynamical systems such as the pharmacokinetic model and wildfire PDE identification, where a batch size of approximately 70-80% of the entire dataset was required, making it nearly equivalent to full-batch optimization. Consequently, only the full-batch ADAM-SINDy results are presented here, as this approach yielded reliable and accurate solutions for all test cases.
@@ 3.3 ABC Flow — Multiple Fourier Frequencies
$\min_{\bm{\Xi}}\left[\|\dot{\mathbf{X}}-\bm{\Theta}(\mathbf{X}\textbf{;}\thinspace\bm{\Lambda})\thinspace\bm{\Xi}\|_{2}^{2}+\lambda\|\hat{\Gamma}\thinspace\bm{\Xi}\|_{1}\right]\;,$ (13)
@@ 3.5 Pharmacokinetics Model — Time-dependent Reaction
where $k_{g}$ denotes the time-dependent reaction rate [63], $k_{0}$ is the initial rate coefficient, and $\eta$ represents the fractal kinetics exponent, constrained within the range $10<\eta<$ 1. The ground truth data are generated using a drug absorption rate of $k_{b}=0.15$, $k_{0}=0.72$, and a fractal kinetics exponent of $\eta=-0.5$, evolved up to $t=10$ with the initial conditions $B_{0}=0$, $G_{0}=1$, and $U_{0}=0$ for a time step of $0.001$.
In this example, the master library is expanded to include $\mathbf{X}\otimes t^{\eta}$, thereby accommodating the time-dependent function, with $\eta$ treated as a trainable parameter within the ADAM-SINDy framework, while a range of parameters $(-0.4,-0.5,-0.55)$ and $(-0.4,-0.55)$ are defined for the SINDy-W and SINDy-W/o methodologies, respectively. Fig. 5 presents a detailed comparison of the governing equations for the ground truth system with those identified by the SINDy-W, SINDy-W/o, and ADAM-SINDy methodologies. Notably, the SINDy-W/o method closely aligns with the ground truth trajectory, exhibiting no deviations. However, the resulting equation from this method is significantly more complex and dense compared to the original system. In contrast, the ADAM-SINDy framework accurately computes the exponent $\eta$ associated with the temporal function, along with other coefficients, mirroring the governing system precisely and thereby demonstrating superior performance compared to the classical SINDy methods.
@@ 3.7 Hyperparameter Robustness — Sparsity Knobs
Our experiments demonstrate that in general the latter optimization strategy, which imposes candidate-wise sparsity, is recommended for nonlinear system identification.
A notable example is the chemical kinetics problem, illustrated in Fig. 7, which represents the former scenario of trainable $\lambda$. It is observed that the SINDy-W method fails to produce the ground-truth equation, whereas ADAM-SINDy remains robust, producing the correct terms with slight deviations in the parameter values when initialized with a higher $\lambda$ value of 0.1. Furthermore, results for the latter scenario with trainable $\bm{\Gamma}$ are not shown here, as the identified terms and parameter values are observed to match the ground truth exactly, as demonstrated in Fig. 4.
@@
$10^{-3}$ $<$ $\lambda$ $<$ $10^{-1}$
$10^{-3}$ $<$ $\lambda$ $<$ 10
$10^{-8}$ $<$ $\bm{\Gamma}_{\sigma}$ $<$ 10
@@
$10^{-3}$ $<$ $\lambda$ $<$ $10^{-2}$
$10^{-5}$ $<$ $\lambda$ $<$ 10
$10^{-5}$ $<$ $\bm{\Gamma}_{\sigma}$ $<$ 10
@@
$10^{-3}$ $<$ $\lambda$ $<$ $10^{-1}$
$10^{-12}$ $<$ $\lambda$ $<$ 10
$10^{-9}$ $<$ $\bm{\Gamma}_{\sigma}$ $<$ 10
@@
$10^{-2}$ $<$ $\lambda$ $<$ $10^{-1}$
$10^{-3}$ $<$ $\lambda$ $<$ 10
1 $<$ $\bm{\Gamma}_{\sigma}$ $<$ 10
@@
$10^{-6}$ $<$ $\lambda$ $<$ 1
$10^{-6}$ $<$ $\bm{\Gamma}_{\sigma}$ $<$ 1

### v3

@@ 1 Introduction
The paper is structured as follows. Section 2 provides an overview of the classical SINDy methodology, outlining its mathematical foundation and then introducing the ADAM-SINDy methodology, emphasizing its advantages and the implementation of the gradient descent algorithm. This section also details the master candidate library of nonlinear functions used in evaluating the classical SINDy and ADAM-SINDy frameworks. Section 3 examines benchmark problems across various applications, including ordinary and partial differential equations, focusing on problems that have nonlinear parameters, thereby challenging the classical SINDy method. Section 4 discusses the advantages of the ADAM-SINDy method over the classical approach. Section 5 presents a conclusion. In the Appendix, we provide more details for the algorithm and provide theoretical and empirical justification for the specific optimization formulation used in ADAM-SINDy.
@@ 2 Methods
2.1 Sparse Identification of Nonlinear Dynamics – SINDy
@@
where $\mathbf{x}(t)\in\mathbb{R}^{n}$ denotes the state vector at time $t$, $\dot{\mathbf{x}}(t)$ represents its time derivative, and $\mathbf{f}(\mathbf{x})$:: $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a nonlinear function governing the system’s dynamics. The objective of the SINDy algorithm is to identify the function $\mathbf{f}(\mathbf{x}(t))$ from available data. The process begins by collecting a dataset consisting of measurements of $m$ states at $n$ discrete time-steps. The data is organized as follows:
@@ 2.2 ADAM Optimization Augmented SINDy: ADAM-SINDy
The ADAM-SINDy methodology, a novel extension of the traditional SINDy framework [34], introduces substantial improvements by adopting ADAM optimization in place of the more commonly employed iterative least squares algorithm for solving the LASSO problem. This methodological innovation significantly addresses key challenges in parameter estimation, while concurrently reducing the dependency on the predefined library of candidate functions, a key limitation that has often constrained the SINDy approaches. Unlike the classical SINDy framework, which heavily relies on the initial selection of candidate parameters, ADAM-SINDy adaptively identifies relevant terms from the candidate library $(\bm{\Xi})$ while simultaneously optimizing nonlinear parameters $(\bm{\Lambda})$. This adaptive process minimizes reliance on prior assumptions and enhances the capability to discover complex, nonlinearly defined parameters in a more robust manner. Nonlinear parameters, such as frequencies in trigonometric functions, bandwidths in exponential functions (e.g., kernels), and non-integer polynomial exponents, present challenges in the traditional SINDy approach, which requires prior knowledge of these parameters. ADAM-SINDy addresses this limitation by making these parameters optimizable within a global optimization framework. The sparse regression problem in Eq. 6 could be reformulated as a min-max or a joint-minimization optimization problem. The min-max optimization problem is written as
$\min_{\bm{\Xi},\bm{\Lambda}}\max_{\lambda,\bm{\Gamma}}\left[\|\dot{\mathbf{X}}-\bm{\Theta}(\mathbf{X};\bm{\Lambda})\bm{\Xi}\|_{2}^{2}+\lambda\|\;|\bm{\Gamma}|\;\bm{\Xi}\|_{1}\right]\;,$ (8)
where the objective is minimized with respect to the coefficient matrix $\bm{\Xi}$ and nonlinear parameters $\bm{\Lambda}$, and maximized with respect to sparsity-controlling variables $\lambda$ and $\bm{\Gamma}$. This min-max approach is similar to the adaptive weighting approach in physics-informed neural networks where the weights for different losses is adaptively updated [57]. Alternatively, the problem can also be expressed as a joint minimization
$\min_{\bm{\Xi},\bm{\Lambda},\lambda,\bm{\Gamma}}\left[\|\dot{\mathbf{X}}-\bm{\Theta}(\mathbf{X};\bm{\Lambda})\bm{\Xi}\|_{2}^{2}+\lambda\|\;|\bm{\Gamma}|\;\bm{\Xi}\|_{1}\right]\;,$ (9)
where all parameters, including coefficients and sparsity weights, are optimized simultaneously by minimizing the objective function. In a first assessment, it appears that the joint minimization leads to $\bm{\Gamma}\rightarrow 0$, effectively removing the $\ell_{1}$ penalty from the loss function and allowing no sparsity in the model output. However, this progression towards zero is counteracted by sequential thresholding applied to both $\bm{\Xi}$ and $\bm{\Lambda}$. Thresholding out candidate functions with negligible contributions forces the corresponding $\bm{\Gamma}$ values to saturate at nonzero levels, while the $\bm{\Gamma}$ associated with significant candidates continues to decrease towards zero. The benefits of this are twofold. First, saturation of $\bm{\Gamma}$ values limits the ability of undesired candidate functions to re-enter the model in later epochs. Second, free movement of $\bm{\Gamma}$ towards zero for remaining candidate terms increases the relative influence of the $\ell_{2}$ data-fidelity loss over the $\ell_{1}$ sparsity penalty, encouraging higher model accuracy as convergence approaches the true candidate functions. The theoretical foundations of this phenomenon, alongside empirical evidence, are discussed further in the B. We utilize joint-minimization for the main results in the paper and discuss min-max in B.
@@
In practice, either $\lambda$ or $\bm{\Gamma}$ is selected to enforce sparsity. Among these, $\bm{\Gamma}$ is preferred due to its ability to treat each candidate function individually. Iteratively updating the weighting matrix $\bm{\Gamma}$ shares similar strategies with the Iteratively Reweighted Least Squares (IRLS) technique, which is applied to both constrained [48] and unconstrained problems [50]. In the present work, we adopt a similar strategy by iteratively updating the $\bm{\Gamma}$ matrix, but with the distinction of updating the weights using gradients computed through backpropagation.
@@
    Adaptive weighting: ADAM-SINDy treats the regularization parameter, $\bm{\Gamma}$, as an optimizable matrix (inspired by the IRLS methodology), where the candidate-specific weights are also optimized over epochs. As a result, the method facilitates an individualized treatment of the candidate functions, thereby enhancing the accuracy of parameter identification and improving the overall robustness of the system identification process. We also show that this approach facilitates achieving machine-precision accuracy.
@@
A detailed description of the ADAM-SINDy algorithm is provided in A, accompanied by the discussion of both stochastic and full-batch optimization strategies. In B, we further provide theoretical/empirical justification for why ADAM-SINDy’s joint-minimization approach together with thresholding works and facilitates machine-precision accuracy.
@@ 2.3 Problem Statement
[…]
@@ 3 Numerical Experiments
Given the two objective formulations considered in this work— joint minimization and min-max optimization—a comparative analysis presented in  B demonstrates that the min-max approach is unsuitable due to the absence of a saddle point. This deficiency hinders convergence and requires further enhancements, such as weight decay, making the approach ultimately less robust. However, theoretical and empirical evidence support the joint minimization formulation as a viable strategy for the identification of nonlinear dynamical systems. Thus, the joint minimization approach has been adopted for the present work. In addition, the results from ADAM-SINDy using batch-wise optimization are excluded, as this method failed to produce accurate outcomes for complex dynamical systems such as the pharmacokinetic model and wildfire PDE identification, where a batch size of approximately 70-80% of the entire dataset was required, making it nearly equivalent to full-batch optimization. Consequently, only the full-batch ADAM-SINDy results are presented here, as this approach yielded reliable and accurate solutions for all test cases.
@@ 3.3 ABC Flow — Multiple Fourier Frequencies
$\min_{\bm{\Xi},\bm{\Lambda}}\left[\|\dot{\mathbf{X}}-\bm{\Theta}(\mathbf{X}\textbf{;}\thinspace\bm{\Lambda})\thinspace\bm{\Xi}\|_{2}^{2}+\lambda\|\hat{\Gamma}\thinspace\bm{\Xi}\|_{1}\right]\;,$ (14)
@@ 3.5 Pharmacokinetics Model — Time-dependent Reaction
where $k_{g}$ denotes the time-dependent reaction rate [63], $k_{0}$ is the initial rate coefficient, and $\eta$ represents the fractal kinetics exponent, constrained within the range $0<\eta<1$. The ground truth data are generated using a drug absorption rate of $k_{b}=0.15$, $k_{0}=0.72$, and a fractal kinetics exponent of $\eta=-0.5$, evolved up to $t=10$ with the initial conditions $B_{0}=0$, $G_{0}=1$, and $U_{0}=0$ for a time step of $0.001$.
In this example, the master library is expanded to include $\mathbf{X}\otimes t^{\eta}$, thereby accommodating the time-dependent function, with $\eta$ treated as a trainable parameter within the ADAM-SINDy framework, while a range of parameters (0.4,0.5,0.55)(-0.4,-0.5,-0.55) and (0.4,0.55)(-0.4,-0.55) are defined for the SINDy-W and SINDy-W/o methodologies, respectively. Fig. 5 presents a detailed comparison of the governing equations for the ground truth system with those identified by the SINDy-W, SINDy-W/o, and ADAM-SINDy methodologies. Notably, the SINDy-W/o method closely aligns with the ground truth trajectory, exhibiting no deviations. However, the resulting equation from this method is significantly more complex and dense compared to the original system. In contrast, the ADAM-SINDy framework accurately computes the exponent $\eta$ associated with the temporal function, along with other coefficients, mirroring the governing system precisely and thereby demonstrating superior performance compared to the classical SINDy methods.
@@ 3.7 Hyperparameter Robustness — Sparsity Knobs
Our experiments demonstrate that the latter optimization strategy, which imposes candidate-wise sparsity, is recommended for nonlinear system identification. A notable example is the chemical kinetics problem, illustrated in Fig. 7, which represents the former scenario of trainable $\lambda$. It is observed that the SINDy-W method fails to produce the ground-truth equation, whereas ADAM-SINDy remains robust, producing the correct terms with slight deviations in the parameter values when initialized with a higher $\lambda$ value of 0.1. Furthermore, results for the latter scenario with trainable $\bm{\Gamma}$ are not shown here, as the identified terms and parameter values are observed to match the ground truth exactly, as demonstrated in Fig. 4.
@@
$10^{-3}$ << $\lambda$ << $10^{-1}$
$10^{-3}$ << $\lambda$ << 10
$10^{-8}$ << $\bm{\Gamma}_{\sigma}$ << 10
@@
$10^{-3}$ << $\lambda$ << $10^{-2}$
$10^{-5}$ << $\lambda$ << 10
$10^{-5}$ << $\bm{\Gamma}_{\sigma}$ << 10
@@
$10^{-3}$ << $\lambda$ << $10^{-1}$
$10^{-12}$ << $\lambda$ << 10
$10^{-9}$ << $\bm{\Gamma}_{\sigma}$ << 10
@@
$10^{-2}$ << $\lambda$ << $10^{-1}$
$10^{-3}$ << $\lambda$ << 10
1 << $\bm{\Gamma}_{\sigma}$ << 10
@@
$10^{-6}$ << $\lambda$ << 1
$10^{-6}$ << $\bm{\Gamma}_{\sigma}$ << 1